OpenStudy (anonymous):

Thirty tickets were sold for a concert, some at 60 cents and the rest at $1. If the total raised was $22, how many had the cheaper tickets?

OpenStudy (anonymous):

let y be the number of tickets sold at $1.00 Eqn 1 x+y=30 Eqn 2 0.6x + 1.0y = 22 y = 30-x substitute y into 2nd equation 0.6x + 1.0(30-x) = 22 0.6x + 30 - x = 22 -0.4x + 30 = 22 0.4x = 8 x = 8/0.4=20 substitute x into eqn 1 20 + y = 30 y = 10